# Improving Zero-noise Extrapolation for Quantum-gate Error Mitigation using a Noise-aware Folding Method

### v1

@@ Abstract
The current thousand-qubit processors mark a substantial advance in hardware. Yet, hardware limitations prevent quantum error correction (QEC), necessitating reliance on quantum error mitigation (QEM). Our paper presents a noise-aware folding method that improves Zero-Noise Extrapolation (ZNE) by estimating noiseless values from noisy results. Unlike traditional ZNE methods, which assume a uniform error distribution, our method redistributes the noise using calibration data based on hardware noise models. By employing noise-adaptive compilation and optimizing the qubit mappings, our approach enhances the ZNE accuracy of various quantum computing models. Recalibrating the noise amplification to address the inherent error variations, promises higher precision and reliability in quantum computations. This paper highlights the uniqueness of our method, summarizes noise accumulation, presents the scaling algorithm, and compares the reliability of our method with those of existing models using linear fit extrapolation. Relative to the existing folding methods, our method achieved a 35% improvement on quantum computer simulators and a 26% improvement on real quantum computers compared to existing folding methods, demonstrating the effectiveness of our proposed approach.
preprint: APS/123-QED
I Introduction
@@
Recently, quantum computing (QC) technology has entered the era of noisy-intermediate scale quantum computing (NISQ), highlighted by IBM’s recent release of processors housing over a thousand qubits [1, 2]; meanwhile, robust crosstalk chips using tunable-coupler technology, can operate up to 133 qubits [3, 4, 5]. Despite these advancements, the capacity of current hardware is inadequate for implementing quantum error correction codes but is projected to become adequate by the 2030s [6, 7]. In the meantime, QC systems employ QEM methods, often requiring additional quantum and classical resources as a trade-off to enhance QC output fidelity.
@@
Numerous QEM techniques for alleviating diverse sources of error in QC have been proposed [8]: probabilistic error cancelation for mitigating decoherence  [9, 10], ZNE for imperfect gates, techniques for mitigating measurement errors  [11, 12, 13], dynamical decoupling [14, 15], quantum optimal control [16, 17], randomized compiling [18, 19], Pauli-frame randomization [20, 21], and other techniques [22, 23, 24, 25, 21, 26]. The present study focuses on the well-established ZNE technique.
@@
ZNE was concurrently introduced and extensively demonstrated in various applications involving systems of up to 127 qubits [27, 28]. In ZNE, a quantum program is amplified to different noise levels through gate or pulse-level stretching, which intentionally increases the program’s noise. Subsequently, the amplified results are extrapolated to estimate the noiseless values. Formally, the program is amplified by multiple scale factors $\lambda$. When $\lambda=1$, the program operates at its original error rate. When $\lambda$ $>0$, an additional error, such as an identity unitary combination of the original gate, is introduced. The results obtained at different noise levels of $\lambda$ are collected and extrapolated to $\lambda$ $=0$, effectively eliminating the noise. Therefore, $\lambda$ represents the noise level affecting any physical quantity during quantum computation.
@@
Researchers have proposed various methods for amplifying noise in quantum programs. For instance, the approach described in [29] stretches gate durations to the desired levels using pulse-level control. Under ideal conditions, this stretching does not alter the quantum system’s state, but under noisy conditions, these modifications are classified as errors. Moreover, as this method requires a high degree of abstract control and calibration at the pulse level in quantum computers, it is not easily implementable across most of the existing QC systems. In constrast, unitary folding [30] amplifies the noise using gate-level control, which is available in all gate-based computation model QCs. This technique builds upon a simple concept: replacing the unitary operation $U\rightarrow U\left(U^{\dagger}U\right)$, where $\left(U^{\dagger}U\right)$ is an identity. In an ideal scenario, this replacement remains a $U$ operation, but errors corresponding to U in the quantum system amplify the errors in the $U$ operation within the modified circuit. Giurgica-Tiron et al.  [30] proposed unitary folding with a fold from the left (which folds each unitary gate independently) and at random folding methods (which randomly selects a subset of individual gates randomly as a block of unitary gates and replicates the entire block with $U\left(U^{\dagger}U\right)$. This method requires no knowledge of the QC underlying noise model. Although these scaling methods effectively amplify noise in the original input quantum program, they neglect the sources of error imbalance within the QC system, potentially leading to biased results and errors that may compromise the accuracy of the extrapolation results.
@@
To resolve this problem, we introduce a noise-aware folding method that adjusts the quantum circuit, aiming to balance the error rate across all logical qubits by leveraging the calibration data noise model, which is freely available on select QC systems. This approach comprehensively analyzes the calibration data noise model, thereby redistributing the noise amplification process across the quantum circuit. Through an algorithmic design, our approach strategically adjusts the scaling factor for each gate operation, balancing the distribution of gate error rates across all logical qubits. Addressing the error source imbalances, mitigates bias-induced errors, thereby enhancing the accuracy of ZNE extrapolated results from the quantum program.
Moreover, the approach is adaptable and therefore versatile across various quantum computing models. Seamlessly integrated with gate-based computation models, it comprehensively caters to the inherent complexities of contemporary quantum hardware. The integration of a noise-aware folding method enhances the precision and reliability of quantum computations. The amplification process can be recalibrated to accommodate the intricate error landscapes within quantum systems, promising more accurate and dependable quantum programming, and bridging the gap between theoretical expectations and practical outcomes. The contributions of our study are highlighted below:
@@
  • Our noise-aware folding method is tailored to redistribute noise across quantum circuits, thereby enhancing the accuracy of ZNE by addressing the inherent error variations within quantum systems.
@@
  • We seamlessly integrate our approach with gate-based computation models and optimize the qubit mappings using hardware noise models, providing versatility mitigating bias-induced errors, and enhancing the reliability of quantum computations.
  • Our novel algorithm dynamically adjusts the error rates within the quantum circuit, leveraging calibration data and noise-adaptive compilation methods to ensure a balanced and controlled scaling process.
@@
  • We conduct rigorous experiments across various full noise models and real quantum computers, showcasing the method’s consistent performance in simulations, highlighting the challenges of scaling to larger circuits, and emphasizing the disparity between simulated and real quantum computer executions.
Figure 1: Controlled-not (CNOT) operation as identity insertion of unitary U1 and U2 operations using two qubits.
@@
The remainder of this is organized as follows. Section II.1 briefly discusses the background and related QEM methods using ZNE techniques, which motivated our research. Section III explains the noise-aware folding method, along with its compilation and execution scheme. Section IV analyzes the performance results of our approach and previous methods. Our noise-aware folding method is further discussed in Section V and conclusions are presented in Section VI.
II Background and Related Works
@@
This section briefly provides the fundamentals of our proposed method. We discuss the quantum error mitigation (QEM), noise scaling methods, and noise extrapolation models.
II.1 Quantum Error Mitigation
Quantum computers are susceptible to various sources of noise and imperfections including crosstalk, which can affect the accuracy, reliability, and fidelity of QC results. Utilizing a set of techniques and strategies to minimize quantum computational errors, QEM estimates the true expectation value $\langle A\rangle$ from the approximated value $A$ impacted by errors $\psi$.
$\langle A\rangle=\operatorname{Tr}(A\psi)$ (1)
QEM strategically minimizes the impact of errors at various stages of quantum processes such as ZNE. Ultimately, it aims to optimize the performance of quantum algorithms and facilitate their seamless integration into practical QC environments. As one of the best-estimated QC techniques, ZNE is used for understanding the behavior of quantum algorithms in the absence of noise. In ZNE, quantum computations are run at different noise levels and the outcomes are extrapolated to a hypothetical noiseless scenario (the zero-noise limit). By analyzing the results obtained at various noise levels, we can predict the performance of a quantum algorithm without noise interference. This extrapolation provides insights into the intrinsic properties and capabilities of the quantum algorithm, clarifying its idealized behavior. Essentially, ZNE can separate the impact of noise from the true potential of a quantum algorithm.
II.2 Noise Scaling Methods
ZNE methods measure a given observable at varying levels of noise. Using the measured dependence on the noise, they then extrapolate the result to the expected noiseless value. Instruction-level noise in current digital quantum computers dominantly arises from the two-qubit entangling controlled-NOT (CNOT) gate, and the dominant noise channel is the two-qubit depolarizing channel [31]. In a two-qubit scenario, the depolarizing channel is given by a quantum operation acting on the system’s density matrix $\rho$:
$\mathcal{E}(\rho)=(1-\epsilon)\rho+\frac{\epsilon}{4}I$ (2)
where $\epsilon$ is the error rate parameter and $I$ is the 2 × 2 identity matrix.
We denote the density matrix of a single noisy (depolarizing) CNOT gate on two-qubit $i$ and $j$ by $\rho_{ij}$ and the unitary operator corresponding to the CNOT gate by $U_{c}$. As the action of two CNOT gates is equivalent to the identity operation (see Figure 1 an odd scaling factor $\lambda$ of CNOT gates must be added to the same qubits as follows
$\operatorname{CNOT}_{ij}^{\lambda}[\rho]=(1-\lambda\epsilon)U_{c}\rho U_{c}+\frac{\lambda\epsilon}{4}\rho_{ij}\otimes I_{ij}+O\left(\epsilon^{2}\right)$ (3)
@@ II.2 Noise Scaling Methods
Unitary folding modifies the quantum circuit by incorporating additional operations or ’folds’ that counteract the effects of noise. This strategically introduces corrective operations that enhance the fault-tolerance of the quantum computation, ultimately mitigating the impact of errors.
Parameterized noise scaling counteracts the effects of noise by adjusting the parameters of quantum gates in a controlled manner to account for and counteract the effects of noise. By dynamically scaling the parameters of quantum gates based on the noise characteristics, this technique optimizes the overall performance of the quantum algorithm in noisy environments.
Several methods for scaling quantum circuits have been reported [30, 29, 32, 33]. The two well-known approaches of Giurgica-Tiron et al. [30], i.e., fold from the left and random folding, manipulate noise solely at the gate 0 level to enable ZNE.
  • Noise Scaling from the Left: This approach selects a specific subset of individual gates or layers in a quantum circuit from the left-hand and subject it to noise scaling operations, starting from the leftmost side of the circuit until the end of the quantum circuit. This approach enables targeted noise scaling of specific portions of the circuit, potentially controlling the overall noise characteristics and error mitigation in a controlled manner.
  • Noise Scaling at Random: This approach randomly selects a subset of individual gates or layers for noise scaling operations. Random gate selection enables uniform sampling of the input circuit, smoothing the convergence to a specific scaling factor. By introducing randomness into the gate selection process, this approach can achieve effective noise scaling and error mitigation in QC.
II.3 Noise Extrapolation Models
Extrapolation models for ZNE in quantum computing have been comprehensively discussed elsewhere [22, 30, 10]. The choice of the model depends on the strength of the noise and the number of gates in the circuit. The four different extrapolation models are summarized below.
  • The linear model [22] assumes that the expectation value of a quantum circuit scales linearly with the number of gates. The linear model is simple and can be useful for estimating the zero-noise limit expected value of the circuit when the noise is weak.
  • The exponential model [22] assumes that the noise in a circuit scales exponentially with the number of gates. This model can be used with unitary folding for exponential scaling of the depolarizing parameters of each gate. The obtained parameters are then used for fitting and extrapolating the expectation value of the circuit.
  • The adaptive-exponential model  [30] an extension of the exponential extrapolation model, allows the scaling factor to vary with the number of gates. Estimating the scaling factor using a Bayesian approach, this method improves the accuracy of the extrapolation and reduces the variance of the estimator.
  • The Richardson model [10] maximizes the order of the polynomial extrapolation to the maximum given the number of data points. When the noise is strong and the polynomial order is high, this model improves the accuracy of the extrapolation and reduces the variance of the estimator.
However, unitary folding [30] might not effectively amplify systematic nonuniform noise on the target quantum hardware. Giurgica-Tiron et al. demonstrated that the adaptive-exponential extrapolation method maximizes the accuracy and minimizes the variance, although the exponential, linear, and Richardson extrapolations are useful in different scenarios.
Figure 2: Overall process of using the noise-aware folding. A traditional qubit mapping process is applied to get a mapped circuit. Using the mapped circuit with the target hardware error rate, we can accumulate the noise on the circuit. Finally, we apply the noise-aware folding method to scale the quantum circuit.
III Proposed Noise-aware Folding ZNE
III.1 Overview

### v2

@@ Abstract
Recent thousand-qubit processors represent a significant hardware advancement, but current limitations prevent effective quantum error correction (QEC), necessitating reliance on quantum error mitigation (QEM) to enhance result fidelity from quantum computers. Our paper introduces a noise-aware folding technique that enhances Zero-Noise Extrapolation (ZNE) by leveraging the noise characteristics of target quantum hardware to fold circuits more efficiently. Unlike traditional ZNE approaches assuming uniform error distribution, our method redistributes noise using calibration data based on hardware noise models. By employing a noise-adaptive compilation method combined with our proposed folding mechanism, we enhance the ZNE accuracy of quantum gate-based computing using superconducting quantum computers. This paper highlights the uniqueness of our method, summarizes noise accumulation, presents the scaling algorithm, and compares the reliability of our method with those of existing models using linear extrapolation model. Experimental results show that compared to existing folding methods, our approach achieved a 35% improvement on quantum computer simulators and a 31% improvement on real quantum computers, demonstrating the effectiveness of our proposed approach.
keywords:
Quantum computing, quantum error mitigation, zero-noise extrapolation, target-aware optimization
1 Introduction
@@
Recently, quantum computing (QC) technology has significant advancements, including IBM’s release of processors housing over a thousand qubits [1, 2], as well as the development of more robust crosstalk chips utilizing tunable-coupler technology with up to 133 qubits [3, 4, 5]. Despite these advancements, the capacity of current hardware is insufficient for implementing QECs which is projected to become available by the 2030s [6, 7]. In the meantime, QC systems employ QEM methods, often requiring additional quantum and classical resources as a trade-off to enhance QC output fidelity.
@@
Numerous QEM techniques for reducing diverse sources of error in QC have been proposed [8]: probabilistic error cancellation for mitigating decoherence  [9, 10], ZNE for imperfect gates, techniques for mitigating measurement errors  [11, 12, 13], dynamical decoupling [14, 15], quantum optimal control [16, 17], randomized compiling [18, 19], Pauli-frame randomization [20, 21], and other techniques [22, 23, 24, 25, 21, 26]. Our study focuses on the well-established ZNE technique.
@@
ZNE was simultaneously introduced and extensively demonstrated in various applications involving systems of up to 127 qubits [27, 28]. In ZNE, a quantum program is amplified to different noise levels through gate or pulse-level stretching, which intentionally increases the program’s noise. Subsequently, the amplified results are extrapolated to estimate the noiseless values. Formally, the program is amplified by multiple scale factors $\lambda$. When $\lambda=1$, the program operates at its original error rate. When $\lambda$ $>1$, additional error, such as an identity unitary combination of the original gate, is introduced. The results obtained at different noise levels of $\lambda$ are collected and extrapolated to $\lambda$ $=0$, effectively reducing noise to improve the result fidelity.
@@
Researchers have proposed various methods for amplifying noise in quantum programs. For instance, the approach described in [29] stretches gate duration to the desired levels using pulse-level control. Under ideal conditions, this stretching does not alter the quantum system’s state, but under noisy conditions, these modifications add additional errors. Moreover, as this method requires a high degree of abstract control and calibration at the pulse level in quantum computers, it is not easily implementable across most of the existing QC systems. In contrast, unitary folding [30] amplifies the noise using gate-level control, which is available in all gate-based QCs. This technique is based on a simple concept unitary folding: replacing the unitary operation $U\rightarrow U\left(U^{\dagger}U\right)$, where $\left(U^{\dagger}U\right)$ is an identity. In an ideal scenario, this replacement remains a $U$ operation, but errors corresponding to U in the quantum system amplify the errors in the $U$ operation within the modified circuit. Giurgica-Tiron et al.  [30] proposed unitary folding with a fold from the left (which folds each unitary gate independently) and with random folding methods (which randomly select a subset of individual gates randomly as a block of unitary gates and replicates the entire block with $U\left(U^{\dagger}U\right)$. This method requires no knowledge of the QC underlying noise model. Although these scaling methods effectively amplify noise in the original input quantum program, they neglect the sources of error imbalance within the QC system, potentially leading to biased results and errors that may compromise the accuracy of the extrapolation results.
@@
To address this issue, we introduce a noise-aware folding method that adjusts the quantum circuit, aiming to balance the error rate across all logical qubits by leveraging the calibration data noise model, which is freely available on most of the QC systems. Our approach comprehensively analyzes the calibration data noise model, strategically adjusts the scaling factor for each gate operation, balancing the distribution of gate error rates across all logical qubits. Addressing the error source imbalances mitigates bias-induced errors, thereby enhancing the accuracy of ZNE extrapolated results from the quantum program.
The contributions of our study are highlighted below.
@@
  • We propose a noise-aware folding method, tailored to redistribute noise across quantum circuits, thereby enhancing the accuracy of ZNE by addressing the inherent error variations within quantum systems.
@@
  • We seamlessly integrate our approach with gate-based computation models and noise-adaptive compilation using hardware noise models, providing versatility, mitigating bias-induced errors, and enhancing the reliability of quantum computations.
@@
  • We conduct rigorous experiments across various full-noise models and real quantum computers, showcasing the method’s consistent performance in simulations, highlighting the challenges of scaling to larger circuits, and emphasizing the disparity between simulated and real quantum computer executions.
@@
The remainder of this paper is organized as follows. Section 2 discusses the background and related QEM methods using ZNE techniques, which motivated our research. Section 3 explains the noise-aware folding method, along with its compilation and execution scheme. Section 4 analyzes the performance results of our approach and previous methods. Analytical results are further discussed in Section 5 and conclusions are presented in Section 6.
2 Background
@@
This section briefly provides the fundamental concept to understand our proposed method. We present QEM, noise extrapolation model, and noise scaling methods.
2.1 Zero-noise Extrapolation Method
Two prominent QEM methods were introduced by Kristan Temme et al. [10] in the same letter. The first method involves extrapolation to the zero-noise limit (known as ZNE), while the second method is known as Probabilistic Error Cancellation (PEC). Both techniques aim to enhance the quality of short-depth quantum circuits and rely on the assumption that a certain noise parameter, combined with system size and circuit depth, can be treated as a small number. In terms of noise tolerance, PEC is more robust than ZNE, but it also requires more resources, which increases exponentially with circuit scale. As a result, ZNE is more suitable for short-term quantum computers, such as those in the NISQ era.
Giurgica-Tiron et al. [30] outlined the fundamental structure of the ZNE problem in their work. The framework assumes that the outcome of a quantum computation is represented by a single expected value denoted as $E(\lambda)$, where $\lambda$ signifies the noise scaling factor. This expected value could emerge from either a solitary quantum circuit or a blend of quantum circuits coupled with classical post-processing. Notably, $E(\lambda)$ is a real number, ideally estimable only under infinite measurement samples. However, in practical scenarios involving a finite number of samples $N$, solely a statistical approximation of the expected value is feasible:
$\hat{E}(\lambda)=E(\lambda)+\hat{\delta}$ (1)
Here, $\hat{\delta}$ stands for a stochastic variable with zero mean and variance $\sigma^{2}$. While $\hat{E}(\lambda)$ is the estimator for $E(\lambda)$, put it simply $\hat{E}(0)$ for $E(\lambda=0)$ is the estimate of the ZNE with its bias and variance. We can draw a genuine prediction $y$ from the probability distribution:
$P(\hat{E}(\lambda)=y)=N(E(\lambda)-y,\sigma^{2})$ (2)
In this distribution, $N(\mu,\sigma^{2})$ represents a typical distribution, often Gaussian, characterized by a mean $\mu$ and variance $\sigma^{2}=\sigma_{0}^{2}/N$. Given a series of $m$ scaling parameters $\lambda=(\lambda_{1},\lambda_{2},...,\lambda_{m})$, where each $\lambda_{j}$ is greater than or equal to 1, and the corresponding outcomes:
$y={y_{1},y_{2},...,y_{m}}$ (3)
Besides the foundational structure ZNE methods, the literature has showcased various models, including linear extrapolation, exponential extrapolation, Richardson extrapolation, among others [30]. In our paper, we will delve into the specifics of the linear extrapolation model, highlighting its utility and presenting experimental results.
Linear extrapolation method stands out as one of the simplest techniques, representing a specific instance within the broader scope of polynomial extrapolation. This method relies on a polynomial model of degree $d$, expressed as:
$E_{poly}^{d}=c_{0}+c_{1}\lambda+...c_{d}\lambda^{d}$ (4)
where $c_{0},c_{1},...,c_{d}$ denote $d+1$ unknown real parameters. From this polynomial model, the linear extrapolation variant emerges:
$E_{linear}(\lambda)=E_{poly}^{(d=1)}(\lambda)=c_{0}+c_{1}\lambda$ (5)
Within this context, a straightforward analytical solution exists, manifesting as the ordinary least squared estimator of the intercept parameter:
$\hat{E}_{linear}(0)=\hat{y}-\frac{S_{\lambda y}}{S_{\lambda\lambda}}\hat{x},$ (6)
$\hat{\lambda}=\frac{1}{m}\sum_{j}{\lambda_{j}},\hskip 42.67912pt\hat{y}=\frac{1}{m}\sum_{j}{y_{j}},$ (7)
[…]
where $\hat{\lambda}$ and $\hat{y}$ represent the averages of the scaling parameters $\lambda_{j}$ and the corresponding outcomes $y_{j}$, respectively. The terms $S_{\lambda y}$ and $S_{\lambda\lambda}$ are computed based on the deviations of $\lambda_{j}$ from their mean $\hat{\lambda}$ and the subsequent squares, facilitating the derivation of the intercept parameter. This analytical solution underscores the efficiency of the linear extrapolation method, offering a streamlined approach to predict outcomes $y$ with respect to input scaling parameters $\lambda$.
2.2 Noise Scaling Methods
Figure 1: Folding methods for ZNE are illustrated with the input circuit depicted in (a), followed by the process of the local unitary folding method to the original circuit in (b), and the global unitary folding method in (c). Noticing the difference in the folding process, where $U^{\dagger}U$ was applied: (b) applies each operation to each of the quantum gates individually, while (c) mirrors the entire circuit as a single $U$ operation.
ZNE method measures a given observable at varying levels of noise. Using the measured dependence on the noise, they then extrapolated the result to the expected noiseless value. We describe the noise scaling method using two main types of folding namely local unitary folding and global unitary folding:
@@ 2.2 Noise Scaling Methods
  • Unitary folding modifies the quantum circuit by incorporating additional operations or ‘folds’, such as global folding and local folding, which counteract the effects of noise [31]. This strategically introduces corrective operations that enhance the fault-tolerance quantum computation, ultimately mitigating the impact of errors.
    $U\rightarrow U\left(U^{\dagger}U\right)^{n},\quad n=0,1,2,\ldots$ (9)
    In the absence of noise, the replacement rule preserves the operation, as $\left(U^{\dagger}U\right)$ equals the identity. However, if noise affects $U$, the unitary folding operation roughly scales the noise by an odd integer factor $\lambda=1+2n$.
  • Global unitary folding: Assume that the circuit is composed of $d$ unitary layers:
    $U=L_{d},...,L_{2},L_{1},$ (10)
    where $d$ represents the depth of the circuit, with each block $L_{j}$ capable of representing a single layer of operations or a solitary gate. Within the global folding approach, the substitution rule from Equation 9 extends across the entire circuit, as depicted in Figure 1(c). This involves duplicating the entire circuit and adding two additional unitary operations $U$, consolidating all gates within. To refine scaling, a final folding can be confined to a subset of the circuit, typically its last $s$ layers. Consequently, the general replacement rule for global folding can be summarized as follows.
    $U\rightarrow U(U^{\dagger}U)^{n}L_{d}^{\dagger}L_{d-1}^{\dagger}...L_{s}^{\dagger}L_{s}...L_{d-1}L_{d}$ (11)
    The total layer count of the modified circuit is $d(2n+1)+2s$, allowing a stretch of the circuit’s depth to be extended up to a resolution scale of $2/d$. This allows for the application of the scaling $d\rightarrow\lambda d$, where
    $\lambda=1+\frac{2k}{d},k=1,2,3,....$ (12)
    Conversely, given any real $\lambda$, one can determine the closest integer $k$ to $d(\lambda-1)/2$, perform integer division of $k$ by $d$ to obtain $n$ and $s$, and then apply $n$ integer foldings along with a final partial folding as detailed in Equation 11.
  • Local unitary folding: Rather than applying global folding to a quantum circuit by appending folds at the end, an alternative approach involves folding a subset of individual gates or layers. As depicted in Figure 1(b), each gate undergoes unitary folding individually using Equation 9. This results in an evident scaling of the initial number of gates or layers, $d$, by an odd integer, $1+2n$. Similar to global folding, a final partial folding operation can be added to achieve a more finely grained scaling factor. To implement such ”partial” folding, let’s define an arbitrary subset, $S$, of the full set of indices ${1,2,...,d}$, where the number of elements in $S$ is a specified integer, $s=|S|$. Within this framework, the following gate folding rule can be defined:
    $\forall_{j}\in\{1,2,...,d\},L_{j}\rightarrow\left\{\begin{array}[]{lr}L_{j}(L_{j}^{\dagger}L_{j})^{n},&if~{}j\notin S\\ L_{j}(L_{j}^{\dagger}L_{j})^{n+1},&if~{}j\in S\end{array}\right.$ (13)
    The selection of elements within the subset $S$ determines the addition of noise channels at various positions throughout the circuit, resulting in various results. In the following section, we will explore the expansion of the ”fold from left” rule into a noise-aware folding mechanism. This approach involves incorporating error rate data to strategically add layers of gates from the left, deviating from the previously proposed method of folding gates a fixed number of times.
Figure 2: The entire process of employing noise-aware folding is described in (a), while (b) illustrates the input data, which represents the target quantum hardware. Our method begins with the application of a traditional qubit mapping process using the input quantum circuit in (a) to generate a mapped circuit for the target hardware topology in (b). Subsequently, the mapped circuit is subjected to the target hardware error rate, which facilitates the accumulation of noise. Following this, the noise-aware folding method is implemented. Initially, the noise on the circuit is accumulated to form an error rate matrix for $\lambda=1$, representing the original mapped circuit. Then, the scaling is applied by computing the scale error rate threshold using the value $\lambda$. Finally, identity gates are inserted until the error rate of the qubit pair approaches the threshold.
3 Proposed Noise-aware Folding ZNE
3.1 Overview
As previously outlined, we adjusted the gate count of the quantum circuit using the scale factor ($\lambda$) through gate-level unitary folding. However, additional techniques such as identity gate insertion [32] and pulse stretching [29] are also utilized to accommodate noise scaling.